3. La Besurta

’PORTILLÓN DE BENÀS’ (OLD MOUNTAIN PORT OF BENÀS) 2 hours

Old mountain passage that used to comunicate the Ball de Benàs with the old French town of Luchón.

Take the asphalt trail that sets off at the end of the main road, which goes to the Llanos del Hospital (Hospital Plains) and the Rencllusa. Once passed the Cabana del Pllan d’Están (Pllan d’Están Hut), you will find a path that goes up the North slope zigzagging to the mountain port. This track is wide and well-traced, and after 1 hour you will reach a landing from where you can already see the opening of the Portillón, on the right (East) of the Salbaguardia Peak. After 2 hours you will get to the border: Portillón de Benàs.

You will enjoy the nice views of the Maladeta and then, crossing a few more metres into France, Boums du Port, you will reach the bottom of the port.